How much work does thr force f =12.0n do on a particule as it moves from x=2.0m to x=7.5


1
Expert's answer
2021-02-22T10:19:42-0500

The work done by the constant force "F", making displacement "s" is "A = F s = F (x_2 - x_1)", where "x_2, x_1" are the final and initial coordinates respectively.

Hence, "A = 12.0 N \cdot (7.5 m - 2.0 m) = 66 J".